OpenStudy (anonymous):

sqaure root of 39 simplified

OpenStudy (anonymous):

\[\sqrt{39} =?\]

OpenStudy (anonymous):

Prime factorize it....

OpenStudy (anonymous):

39 = 3*13.....it can be further simplified......but u can find ita value \[\sqrt{39}=6.24\]

OpenStudy (anonymous):

calculating a numeric value of a square root isn't usually what they mean by "simplify".

OpenStudy (anonymous):

the hyptenuse of a right triangle is 8cm. If one leg of the triangle is 5cm. Find the other leg. Simplify your answer. <<< this is the question

OpenStudy (anonymous):

ok, great... a^2 + b^c = c^2 5^2 + b^2 = 8^2 25 + b^2 = 64 b^2 = 64 - 25 = 39 b = sqrt(39) That's all you can do... there's no way to simplify sqrt(39). If they want you to have a numerical answer, you can use a calculator, but typically, you just leave it as sqrt(39)

OpenStudy (anonymous):

i meant hypotenuse

OpenStudy (anonymous):

usually, "simplify" a sqrt means something like "simplify sqrt(50)" and you are supposed to realize that the stuff under the square root includes a factor that is a perfect square, like in this case, 50 = (25)(2) So to simplify sqrt(50), you get sqrt(25 x 2) = 5sqrt(2)

OpenStudy (anonymous):

lol....@JakeV8 thats wat i told...)

OpenStudy (anonymous):

I know you did, @Yahoo! ... and you were right, too... @BrierraT just seemed to need a little more explanation.

OpenStudy (anonymous):

thanks !